OpenStudy (anonymous):

Can someone walk me through these and help me get the right answer .... A quadrilateral has vertices (2, 0), (0, –2), (–2, 4), and (–4, 2). Which special quadrilateral is formed by connecting the midpoints of the sides? kite rectangle trapezoid rhombus 2. Which of the following describes TVS? The vertices are T(1, 1), V(4, 0), and S(3, 5) isosceles scalene right equilateral 4. Verify that parallelogram ABCD with vertices A(–5, –1), B(–9, 6), C(–1, 5), and D(3, –2) is a rhombus by showing that it is a parallelogram with perpendicular diagonal

OpenStudy (anonymous):

Use a graph to solve 1

OpenStudy (anonymous):

er solve them all

OpenStudy (angeleyes108):

This is biology. You might want to post this in either the math category or the Algebra category.

OpenStudy (anonymous):

Oh lol , I didnt even realize I was in the bio section. Thanks For letting me know !!

OpenStudy (angeleyes108):

That's okay, Ive done it before too lol

OpenStudy (angeleyes108):

But I do know some of these if you need help. 1. rectangle.

OpenStudy (angeleyes108):

2. scalene

OpenStudy (angeleyes108):

And 4. Determine the two sets of opposite edges are parallel to each other: slope of AB = difference of y's / difference of x's = (6-(-1))/(-9- (-5)) = 7/(-4) slope of CD = (-2-5)/(3-(-1)) = -7/4 so AB || CD similarly, slope of BC = (5-6)/(-1-(-9)) = -1/8 slope of DA = (-1-(-2))/(-5-3) = 1/(-8) so BC || DA

OpenStudy (angeleyes108):

The diagonals: slope of AC: (5-(-1))/(-1-(-5)) = 6/4 = 3/2 slope of BD: (-2-6)/(3-(-9)) = -8/12 = -2/3 since the slopes are negative reciprocals of each other, they are perpendicular. The unspoken assumption is that usually, the vertices are numbered in order as you go around the shape, not left to right in rows like you're reading.

OpenStudy (angeleyes108):

Hope i helped. :)

OpenStudy (anonymous):

:(